OpenStudy (anonymous):

find the parameterization of the semi-circle given by x^2+y^2=255, y<=0 in the xy-plane starting at the point (-15, 0) and ending at (15, 0)

OpenStudy (anonymous):

You mean writing it down as a parametric equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

parametric equation of circle having the origin as centre is usually in the form x = a cos t, y = a sin t, where a is the radius of the circle.

OpenStudy (anonymous):

\[(15 \cos t, 15 \sin t) \qquad0\le t \le \pi\]

OpenStudy (anonymous):

to be honest, i dont even think this was covered in class but it's in HW, not sure if the professor made a mistake, and i dont even know how the answer should look like, btw this is vector cal, so everything should be in 3d

OpenStudy (anonymous):

Yeah, that's right xactxx